OpenStudy (anonymous):

If P(x+1)=3x^2-2x-5, Find P(4).

OpenStudy (anonymous):

@waterineyes may answer is P= 12x-20 am I right?

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

u should get a full value with no variables. since, x + 1 = 4 hence, x = 3. sub this in the eq.

OpenStudy (anonymous):

Thanks alot

OpenStudy (anonymous):

In place of x you have to substitute 3: See, P(x+1) Here we have to find P(4) so, x + 1 = 4 or x = 3 Now put this x = 3 in the given function: \[3(3)^2 - 2(3) - 5\] Calculate its values..

OpenStudy (anonymous):

16

OpenStudy (anonymous):

Yes you are absolutely right..

OpenStudy (anonymous):

what i did was i factor it so I get 12x-20

OpenStudy (anonymous):

After factoring you can put the value also: 12x - 20 = 12*3 - 20 = 36 - 20 = 16.. The main thing is to put the values of x here..

OpenStudy (anonymous):

Thanks a lot guys!

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

@waterineyes can you please help me in my attached files. My teacher said it has an answer but my mom said It doesn't have

OpenStudy (anonymous):

My internet speed is slow so it is taking time in opening so wait please.

OpenStudy (anonymous):

it is ok dear

OpenStudy (anonymous):

Line I is parallel to what?? And Line II is perpendicular to what??

OpenStudy (anonymous):

That is what I told to my teacher too and he send me this email Problem 1 (a) is not as confusing as you think. All you have to do is rearrange both equations to solve for y in each case and see what the slopes, the coefficient of x are. You can then figure out what A1/B1 is in terms of A2 and B2, if the lines are assumed to be parallel or perpendicular to each other. @waterineyes

OpenStudy (anonymous):

Yes I got it now: Write here your first equation..

OpenStudy (anonymous):

It is in attached file cause dont know how to use drawing here

OpenStudy (anonymous):

Okay I will do it for you..

OpenStudy (anonymous):

Thanks a lot. This is in my review paper for finals.Thanks a lot @waterineyes

OpenStudy (anonymous):

\[A_1x + B_1y = C_1\] Evaluate y from here: \[y = \frac{C_1 - A_1x}{B_1}\] Or we can write it as: \[y = \frac{-A_1}{B_1}x + \frac{C_1}{B_1}\] Now can you tell me coefficient of x here??

OpenStudy (anonymous):

I dont understand still

OpenStudy (anonymous):

I show you step by step but do you know about coefficient??

OpenStudy (anonymous):

is a multiplicative factor in some term of an expression @waterineyes

OpenStudy (anonymous):

Yes.. Suppose you are given with -5x.. So, -5 is the coefficient here..

OpenStudy (anonymous):

is it -A,B and C?

OpenStudy (anonymous):

I will come to it later.. See, you are given with: \(A_1x + B_1y = C_1\) Subtract both the sides by \(A_1x\) You will get: \[B_1y = C_1 - A_1x\] For simplicity write the x terms in front: So, \[B_1y = -A_1x + C_1x\] Divide by \(B_1\) both the sides: \[y = \frac{-A_1}{B_1} + \frac{C_1}{B_1}\] Till here any difficulty??

OpenStudy (anonymous):

There will not come \(C_1x\) it is simply \(C_1\) in the second last step..

OpenStudy (anonymous):

so this is the answer for parallel in what number and perpendicular?

OpenStudy (anonymous):

This is only 25 % complete..

OpenStudy (anonymous):

so there is a lot of steps to get the answers?

OpenStudy (anonymous):

Have you got the steps tell me honestly...

OpenStudy (anonymous):

the one you teach me yes

OpenStudy (anonymous):

So, we got: \[y = \frac{-A_1}{B_1}x + \frac{C_1}{B_1}\] Now tell me what is the term attached to x here??

OpenStudy (anonymous):

the A2?

OpenStudy (anonymous):

In the equation I have just written above.. i think there is \(\frac{-A}{B}\) attached to the x.. right??

OpenStudy (anonymous):

Sorry.. \[\frac{-A_1}{B_1}\]

OpenStudy (anonymous):

i dont know.I just understand it when you try to explain it to me

OpenStudy (anonymous):

yes it is attached

OpenStudy (anonymous):

That is what I am trying to say.. This is called the coefficient of x.. So, this is called slope also in case of Lines.. It is another vast concept.. But here you can remember that the term attached to x is called slope if the given line is of the form: y = mx + c Here m is the slope.. So, \[Slope = \frac{-A_1}{B_1}\] Okay??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

therefore?

OpenStudy (anonymous):

Similarly find the Slope of second given equation: \[A_2x + B_2y = C_2\] I am writing directly: \[y = \frac{-A_2}{B_2}x + \frac{C_2}{B_1}\] So, slope of the second line is : \[Slope_2 = \frac{-A_2}{B_2}\] Got till here??

OpenStudy (anonymous):

@waterineyes this is parallel to A1+B1=C1

OpenStudy (anonymous):

A1x+B1y=C1

OpenStudy (anonymous):

I am coming to all your doubts.. Firstly answer me do you get these steps also?? Till now we have calculated the slope of first and second line..

OpenStudy (anonymous):

yes, You explain it good to me steb by steb

OpenStudy (anonymous):

So, now I am writing it in CAPITAL: IF THE TWO LINES ARE PARALLEL, THEN THERE SLOPES ARE EQUAL.. Can you remember this for lifetime??

OpenStudy (anonymous):

yes and if it is perpendicular lines consist of two slopes are negative reciprocal

OpenStudy (anonymous):

I remembered my teacher teach us this

OpenStudy (anonymous):

So, according to this property: \[\color{green}{\frac{-A_1}{B_1} = \frac{-A_2}{B_2}}\] Finally we have found the relation in case of PARALLEL LINES: \[\large \color {blue}{\frac{A_1}{B_1} = \frac{A_2}{B_2}}\]

OpenStudy (anonymous):

Thanks a lot

OpenStudy (anonymous):

Hawwwwwwww.. You have been taught this thing..... Now this becomes easy then.. Now, in Case of Perpendicular, one slope is negative reciprocal of the other. So, in case of PERPENDICULAR LINES: \[\large \color{green}{\frac{-A_1}{B_1} = -(\frac{B_2}{-A_2})}\] Or, \[\large \color{blue}{\frac{A_1}{B_1} = \frac{-B_2}{A_2}}\]

OpenStudy (anonymous):

so this is the perpendicular?

OpenStudy (anonymous):

yeah he is explain it a little in our class though

OpenStudy (anonymous):

@waterineyes are u a teacher? U teach us good

OpenStudy (anonymous):

No, I am still a student.. I Love Mathematics.. In School Times, my Math Teacher Mr. S Arjunan was an excellent teacher.. Today, What I know about Mathematics is just because of him..

OpenStudy (anonymous):

Instead I am in College now..

OpenStudy (anonymous):

@waterineyes where u from? He did well in teaching you. You are definitely Great!

OpenStudy (anonymous):

I am from \(\large \color{green}{INDIA}..\)

OpenStudy (anonymous):

my Indian Classmates are smart in Math too.

OpenStudy (anonymous):

Oh that is the really nice thing..

OpenStudy (anonymous):

@waterineyes we are done right? my teacher will be herein a minute.

OpenStudy (anonymous):

They are excellent in terms of Math here.

OpenStudy (anonymous):

Yes we have done..